On October 31, Legacy Rocks Inc., a marble contractor, issued for cash 100,000 shares of $8 par common stock at $10, and on November 19, it issued for cash 50,000 shares of $15 par preferred stock at $51. a. Journalize the entries for October 31 and November 19. For a compound transaction, if an amount box does not require an entry, leave it blank. Oct. 31 Nov. 19 b. What is the total amount invested (total paid-in capital) by all stockholders as of November 19? $
Answer #1
a
Oct .31 Cash 1000000 =100000*10
       Common Stock 800000 =100000*8
       Paid in Capital in excess of par-Common Stock 200000
Nov. 19 Cash 2550000 =50000*51
       Preferred Stock 750000 =50000*15
       Paid in Capital in excess of par-Preferred Stock 1800000
b
Total amount invested (total paid-in capital) 3550000 =1000000+2550000